# Optimal Allocations under Strongly Pigou-Dalton Criteria: Hidden Layer Structure & Efficient Combinatorial Approach

### v1

@@ 4.2. Find a stable allocation for divisible items using LP and a proof of Theorem 17
Finally, recall the result of Nace and Doan (2002) which states that the fair multi-flow problem can be reduced to a polynomial number of linear programming problems and thus be solved in polynomial time.
Proof of Theorem 17.
Recall finding a stable allocation by computing a fair multi-flow, the flow vector $(|f_{1}|,\ldots,|f_{n}|$) is unique for multi flow problem, according to Nace and Doan (2002). Therefore the profile of optimal allocation is also unique. ∎
5. Indivisible items and agents with 0/1-sub valuations
@@
The 0/1-sub function is closely related to the matroid theory Oxley (2011). A matroid is a pair $(E,\mathcal{I})$, where $E$ is a finite set (called the ground set) and $\mathcal{I}$ is a family of subsets of $E$ (called the independent sets).
@@
If an agent has a 0/1-sub valuation on items, then the set of clean bundles forms the set of independent sets of a matroid (proved in Benabbou et al. Benabbou et al. (2021)). Benabbou et al. Benabbou et al. (2021) further showed that under 0/1-sub valuations, the set of max-USW allocations (under any SPD criterion) are consistent with $\mathsf{LexiMin}$ allocations. This result generalizes Theorem 7 Claim 1 (for the 0,1-add valuations). Besides, Babaioff et al.Babaioff et al. (2021) showed that under 0/1-sub valuation, the allocation that optimizes $\mathsf{NSW}^{\neg}$ can be found in polynomial time.
@@
Recall that the family of clean bundles $\mathcal{I}_{j}=\{S\subseteq[m]\mid v_{j}(S)=|S|\}$ for $j\in[n]$ forms a family of independent sets of a matroid Benabbou et al. (2021). By ((I3)) of the independent-set matroid axioms and inequality (7), there exists an item $o_{1}\in\chi^{\prime}_{i}\setminus\chi_{i}$ making $v_{i}(\chi_{i}\cup\{o_{1}\})=v_{i}(\chi_{i})+1=h_{i}+1$. And $o_{1}$ is allocated to some agent $i_{1}\neq i$ under allocation $\chi$. Otherwise, $o_{1}$ is not allocated to anyone, and can be allocated to agent $i$ violating that $\chi$ is max-USW. Consider the following three cases:
  1. (1)
@@
Figure 3. Transferring items along the sequence $(i_{0},o_{1},i_{1},\dots,o_{t},i_{t})$. Some items are denoted by $a,\dots,g$.
To sum up, suppose to the opposite of Theorem 22 that there are some pairs of allocations $(\chi,\chi^{\prime})$ that meets (i) $\chi,\chi^{\prime}\in\mathcal{SS}$ and (ii) $D(\mathbf{p}(\chi),\mathbf{p}(\chi^{\prime}))\geq 2$, then it leads to some contradiction. Thus, for $\chi,\chi^{\prime}\in\mathcal{SS}$, it holds that $D(\mathbf{p}(\chi),\mathbf{p}(\chi^{\prime}))\leq 1$. ∎
Remark 0.
Our proof of Theorem 22 is similar to that of Lemma 3.12 in Benabbou et al. (2021), which is the 0/1-sub valuation version of Lemma 6 (implying an allocation non-optimal under $\mathsf{LexiMin}$ is non-optimal under any SPD criterion). There is a minor flaw in their proof of Lemma 3.12 in Benabbou et al. (2021) which we have modified in our proof of Theorem 22.
In the proof of Lemma 3.12 in Benabbou et al. (2021), for the sequence of items and agents $(i_{0},o_{1},i_{1},\dots,o_{t},i_{t})$, they argue that no same agent appears twice, which implies that if the same agent appears again, the allocation $\chi$ is still clean after transferring items along the cycle (which is a subsequence of the sequence). However, this is not true. See the example in Figure 3, where the sequence is $(3,a,2,b,1,c,2,d,4)$. After transferring items along the cycle $(2,b,1,c,2)$ (which is equivalent to swapping items $b$ and $c$ between agents $2$ and $1$), the allocation may not be clean if $\{a,b\}$ is not a clean bundle of agent $2$. Indeed, the family of clean bundles of agent $2$ may be $\mathcal{I}_{2}=\{\{a,c\},\{a,d\},\{b,c\},\{b,d\},\{a\},\{b\},\{c\},\{d\}\}$ such that $\{a,b\}\notin\mathcal{I}_{2}$.
We note that Lemma 3.12 in Benabbou et al. (2021) still holds and the minor flaw in its proof in Benabbou et al. (2021) can be corrected by arguing that no same item appears twice like ours proof of Theorem 22.
@@
Appendix A All Criteria in Example 2 are SPD
Claim 1.
@@ Appendix A All Criteria in Example 2 are SPD
    (Hence $\mathsf{GiniIndex}(\mathbf{q})<\mathsf{GiniIndex}(\mathbf{p})$ due to 1.)
Therefore all criteria mentioned in Example 2 are SPD.
Proof.
@@
Thus we have $\mathsf{EnvySum}(\mathbf{q})-\mathsf{EnvySum}(\mathbf{p})<0$. ∎
Appendix B Proofs of Lemma 6 and Theorem 9
Proof of Lemma 6.
Note that $\chi^{(b)}$ becomes $\chi^{*}$ after applying the arcs in $C_{1},\ldots,C_{a}$. We obtain that $\mathsf{LexiMin}(\chi^{(b)})=\mathsf{LexiMin}(\chi^{*})$. Further since that $\mathsf{LexiMin}(\chi^{*})<\mathsf{LexiMin}(\chi)$, there exists $i\penalty 10000\ (1\leq i\leq b)$ such that $\mathsf{LexiMin}(\chi^{(i)})<\mathsf{LexiMin}(\chi^{(i-1)})$.
It follows that in $\chi^{(i-1)}$, we have $h_{s}\geq h_{t}+2$ (where $s,t$ denote $s_{i},t_{i}$ respectively, for short). It further follows that in $\chi^{(0)}=\chi$, we also have $h_{s}\geq h_{t}+2$, as $h_{s}$ never increases and $h_{t}$ never decreases in the sequence $\chi^{(0)},\ldots,\chi^{(i-1)}$. Consequently, $\chi$ admits a narrowing transfer (from $s$ to $t$). ∎
Proof of Theorem 9.
Decompose $G$ into paths $P_{1},\ldots,P_{b}$ and a few cycles. Denote by $s_{j},t_{j}$ the starting and ending vertices of $P_{j}$, respectively. We assume that $t_{j}\neq s_{k}$ for $j\neq k$ as in Lemma 6. Moreover, assume that $s_{1}=s_{2}=i$ (due to $h_{i}\geq h^{\prime}_{i}+2$).
Because $\mathbf{p}(\chi^{(b)})=\mathbf{p}(\chi^{\prime})$, we have $\mathsf{LexiMin}(\chi^{b})=\mathsf{LexiMin}(\chi^{\prime})$. We also know $\mathsf{LexiMin}(\chi^{(0)})=\mathsf{LexiMin}(\chi^{\prime})$ since both $\chi^{(0)},\chi^{\prime}$ are stable (Lemma 6). Together, $\mathsf{LexiMin}(\chi^{(0)})=\mathsf{LexiMin}(\chi^{(b)})$. It further implies that $\mathsf{LexiMin}(\chi^{(0)})=\mathsf{LexiMin}(\chi^{(1)})=\ldots=\mathsf{LexiMin}(\chi^{b})$. Otherwise, there exists $j$ such that $\mathsf{LexiMin}(\chi^{(j)})$ ¡ $\mathsf{LexiMin}(\chi^{(j-1)})$, which means $\chi^{(j-1)}$ admits a narrowing transfer $s_{j}\rightarrow t_{j}$, implying that $\chi^{(0)}$ admits a narrowing transfer $s_{j}\rightarrow t_{j}$.
Since $\mathsf{LexiMin}(\chi^{(0)})=\mathsf{LexiMin}(\chi^{(1)})=\mathsf{LexiMin}(\chi^{(2)})$, we know $s_{1}\rightarrow t_{1}$, i.e. $i\rightarrow t_{1}$ (recall $s_{1}=i$), is a swapping transfer of $\chi^{(0)}$, whereas $s_{2}\rightarrow t_{2}$, i.e. $i\rightarrow t_{2}$ (recall $s_{2}=i$), is a swapping transfer of $\chi^{(1)}$. It follows that $h_{t_{2}}=h_{i}-2$, therefore $\chi^{(0)}=\chi$ admits a narrowing transfer $i\rightarrow t_{2}$. ∎
Appendix C Find a stable allocation for indivisible items in $O(m^{2}n)$ time
According to Theorem 7 Claim 1, finding a stable allocation reduces to finding the allocation with minimum $\mathsf{Congestion}$.
Obviously, the latter further reduces to computing the minimum-cost flow in the following network (see Figure 4):
Figure 4. Reduction to minimum-cost flow problem.
Our target – a flow of size $m$ with the minimum cost – can be computed by the Successive Shortest Path algorithm, which increases the size of the current flow by 1 via augmenting along the shortest path in the residual graph, repeating $m$ times. For our particular network, finding such a path reduces to finding a non-used edge $(v_{j},t)$ with lowest cost such that $s$ can reach $v_{j}$ in the residual network, which can be done in $O(mn)$ time by BFS. In total it runs in $O(m^{2}n)$ time.
Appendix D Inconsistency for the Mixed Case

### v3

@@ 4 Divisible items and agents with 0/1-add valuations
Finally, recall the result of [nace2002polynomial] which states that the fair multi-flow problem can be reduced to a polynomial number of linear programming problems and thus be solved in polynomial time.
Proof(Proof of Theorem˜4.2)
Recall finding a stable allocation by computing a fair multi-flow, the flow vector $(|f_{1}|,\ldots,|f_{n}|$) is unique for multi flow problem, according to [nace2002polynomial]. Therefore the profile of optimal allocation is also unique.
5 Indivisible items and agents with 0/1-sub valuations
@@
The 0/1-sub function is closely related to the matroid theory [Oxley-10.1093/acprof:oso/9780198566946.001.0001]. A matroid is a pair $(E,\mathcal{I})$, where $E$ is a finite set (called the ground set) and $\mathcal{I}$ is a family of subsets of $E$ (called the independent sets).
@@
If an agent has a 0/1-sub valuation on items, then the set of clean bundles forms the set of independent sets of a matroid (proved in Benabbou et al. [10.1145/3485006]). Benabbou et al. [10.1145/3485006] further showed that under 0/1-sub valuations, the set of max-USW allocations (under any SPD criterion) are consistent with $\mathsf{LexiMin}$ allocations. This result generalizes Theorem˜3.1 Claim 1 (for the 0,1-add valuations). Besides, Babaioff et al.[babaioff2021fair] showed that under 0/1-sub valuation, the allocation that optimizes $\mathsf{NSW}^{\neg}$ can be found in polynomial time.
@@
Recall that the family of clean bundles $\mathcal{I}_{j}=\{S\subseteq[m]\mid v_{j}(S)=|S|\}$ for $j\in[n]$ forms a family of independent sets of a matroid [10.1145/3485006]. By ((I3)) of the independent-set matroid axioms and inequality (5), there exists an item $o_{1}\in\chi^{\prime}_{i}\setminus\chi_{i}$ making $v_{i}(\chi_{i}\cup\{o_{1}\})=v_{i}(\chi_{i})+1=h_{i}+1$. And $o_{1}$ is allocated to some agent $i_{1}\neq i$ under allocation $\chi$. Otherwise, $o_{1}$ is not allocated to anyone, and can be allocated to agent $i$ violating that $\chi$ is max-USW. Consider the following three cases:
  1. 1.
@@
Figure 4: Transferring items along the sequence $(i_{0},o_{1},i_{1},\dots,o_{t},i_{t})$. Some items are denoted by $a,\dots,g$.
To sum up, suppose to the opposite of Theorem˜5.1 that there are some pairs of allocations $(\chi,\chi^{\prime})$ that meets (i) $\chi,\chi^{\prime}\in\mathcal{SS}$ and (ii) $D(\mathbf{p}(\chi),\mathbf{p}(\chi^{\prime}))\geq 2$, then it leads to some contradiction. Thus, for $\chi,\chi^{\prime}\in\mathcal{SS}$, it holds that $D(\mathbf{p}(\chi),\mathbf{p}(\chi^{\prime}))\leq 1$.
Remark 5
Our proof of Theorem˜5.1 is similar to that of Lemma 3.12 in [10.1145/3485006], which is the 0/1-sub valuation version of Lemma˜1 (implying an allocation non-optimal under $\mathsf{LexiMin}$ is non-optimal under any SPD criterion). There is a minor flaw in their proof of Lemma 3.12 in [10.1145/3485006] which we have modified in our proof of Theorem˜5.1.
In the proof of Lemma 3.12 in [10.1145/3485006], for the sequence of items and agents $(i_{0},o_{1},i_{1},\dots,o_{t},i_{t})$, they argue that no same agent appears twice, which implies that if the same agent appears again, the allocation $\chi$ is still clean after transferring items along the cycle (which is a subsequence of the sequence). However, this is not true. See the example in Figure˜4, where the sequence is $(3,a,2,b,1,c,2,d,4)$. After transferring items along the cycle $(2,b,1,c,2)$ (which is equivalent to swapping items $b$ and $c$ between agents $2$ and $1$), the allocation may not be clean if $\{a,b\}$ is not a clean bundle of agent $2$. Indeed, the family of clean bundles of agent $2$ may be $\mathcal{I}_{2}=\{\{a,c\},\{a,d\},\{b,c\},\{b,d\},\{a\},\{b\},\{c\},\{d\}\}$ such that $\{a,b\}\notin\mathcal{I}_{2}$.
We note that Lemma 3.12 in [10.1145/3485006] still holds and the minor flaw in its proof in [10.1145/3485006] can be corrected by arguing that no same item appears twice like ours proof of Theorem˜5.1.
@@
Appendix 0.A All Criteria in Example 1 are SPD
Claim
@@ Appendix 0.A All Criteria in Example 1 are SPD
    (Hence $\mathsf{GiniIndex}(\mathbf{q})<\mathsf{GiniIndex}(\mathbf{p})$ due to Claim 0.A.)
Therefore all criteria mentioned in Example˜1 are SPD.
Proof
@@
Thus we have $\mathsf{EnvySum}(\mathbf{q})-\mathsf{EnvySum}(\mathbf{p})<0$.
Appendix 0.B Inconsistency for the Mixed Case